OpenStudy (anonymous):

A refrigerator has a width of 3 feet. The angle of elevation from the bottom corner to the opposite corner is 65º. What is the height of the refrigerator?

OpenStudy (anonymous):

A. 6.4 feet B. 7 feet C. 7.4 feet

OpenStudy (anonymous):

Is the depth also 3 feet?

OpenStudy (amistre64):

tan(h/3) = 65 h/3 = tan-1(65) h = 3 tan-1(65)

OpenStudy (amistre64):

err.. got my inverses backwards :) just tan it

OpenStudy (amistre64):

h = 3 tan(65)